OpenStudy (anonymous):

Help please idk if I'm getting the right answer 5x - 8(x + 2) = -3(x - 4)

OpenStudy (anonymous):



OpenStudy (gabylovesyou):

-3x - 16 = 12 - 3x

OpenStudy (gabylovesyou):

now: Whats next?

OpenStudy (gabylovesyou):

add 3x to both sides, what happens?

OpenStudy (gabylovesyou):

x is some number so if x+2 = x+4 then subtracting x from both sides gives us 2 = 4 which is always false so there are no solutions to x+2 = x+4

OpenStudy (gabylovesyou):

so -3x+something = -3x+something else is the same kind of contradiction -----> no solution